OpenStudy (anonymous):

The product of 14 and n is -28. Write the algebraic equation A number decreased by 32 is -58. Find the number.

OpenStudy (anonymous):

14 * n = -28 n - 32 = -58, n = -58 +32 = -26

OpenStudy (anonymous):

thnkss

OpenStudy (anonymous):

You're welcome.